OpenStudy (anonymous):

NEED HELP ASAP ! Factor 6x^3+5x^2-18x-15

OpenStudy (roadjester):

This type of problem is mostly trial and error. What you want to focus on is the first and last term. You have a \(\large6x^3\) and a -15. From that you want to cross-multiply the terms until one gets you 5 and one gets you -18.

OpenStudy (roadjester):

Now, \(\large +5x^2\) can only be generated with the 5 coming from the 15.

OpenStudy (roadjester):

That tells you that you're going to be using 3 and 5 in your constants with one being positive and one negative.

OpenStudy (roadjester):

Ok, fine just give the answer away. @eliassaab =(

OpenStudy (anonymous):

People should start using wolframalpha for this kind of questions and try to figure out themselves how wolframalpha got the answer.

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

YW

OpenStudy (roadjester):

Fine, just pretend like I don't exist. =(

OpenStudy (anonymous):

Here how you do it by hand \[ 6x^3+5x^2-18x-15= 6x^3-18x+5x^2-15=\\ 6x( x^2 -3) + 5(x^2-3)=(6x +5)(x^2-3)=(6x+5)( x- \sqrt 3)(x+\sqrt 3) \]